### v1

@@ 3.1.2 The Lieb-Liniger one dimensional model for interacting bosons
with $x_{j}$ being the position of the $j$-th particle on the ring and $\phi_{0}$ corresponding to the bare flux quantum. The solution of this model provides us with a full many-body tractable model of interacting bosons in a finite-size ring that can sustain persistent currents. The Lieb-Liniger model is integrable through Bethe-ansatz [220] and will be discussed in detail in the following section. Its integrability and exact solution allow us to use this model as a test bench and reference model to understand the microscopic origin of the phenomena observed in more complex systems. For this reason, a plethora of works have and still do investigate bosonic systems using the Lieb-Liniger model [221, 222, 223, 119]. Despite its integrability through Bethe Ansatz, calculating high-order correlations using the exact solution beomes difficult. New techniques have been proposed, for example the use of quantum computers, in which one can prepare the quantum state to later perform measures on it [224, 225].
The two models presented, GPE and LL, describe bosonic interacting particles in the continuum. However we remark again that the first is based on mean-field theory and therefore restricted to the limit of weak interactions and a large number of particles, while the latter describes the full-many body wavefunction exactly. Both theories can attain attractive, $a_{s}<0$, and repulsive $a_{s}>0$ interactions. For repulsive interactions, both models give a ground state with a homogeneous density along the ring because of the translational invariance present in the system. Intuitively, repulsion tends to spread particles apart, and therefore, a homogenous density keeps them as separated as possible without creating regions of higher density that would lead to increased local interactions. On the other hand, for attractive interactions, one could expect the localization of particles in space due to the negative two-body contact interactions. Such localization would, in principle, break translational invariance, and it is found in the mean-field description for sufficiently strong attractive interactions [226]. However, the apparent breaking of translational invariance is only observed in single-shot density measurements where the particle’s center-of-mass is localized at a given position. One can also access such localization by calculating expectation values of higher order correlators. We remark that, as pointed out in [221], the alleged breaking of translational invariance is because one would need to average over the center-of-mass position of the mean-field solution in order to recover the correct translational invariant solution. This connection between mean-field and many-body theories were also investigated in Ref. [227].
@@ 3.1.3 The Bose-Hubbard model
where $b^{\dagger}_{j}$ ($b_{j}$) creates (annihilates) a boson, and $n_{j}=b_{j}^{\dagger}b_{j}$ is the local particle number operator for site $j$. The parameters $J$ and $U$ account for the hopping amplitude and on-site interaction when restricting the particles to occupy the lowest Bloch band in each lattice site respectively [228]. The elementary flux quantum is denoted by $\phi_{0}$ and is taken to be 1 unless explicitly stated. The BHM has been experimentally realized in experiments [63, 65, 229] with remarkable control on the parameters of the systems and read-out of the particles trapped in the lattice. With such a model, they probe the superfluid to Mott insulator transition in a highly controllable environment. This model also provides a full many-body description of the interacting bosons. However, due to its imposed restriction provided by the lattice, the Hilbert space dimensions become approachable from a numerical perspective, with techniques such as Exact Diagonalization or density matrix renormalization group (DMRG) [230, 231].
@@ 3.1.5 Exact solutions for interacting bosons in 1D: Bethe Ansatz and Tonks-Girardeau
Repeated use of Eq. (36) together with the imposition of the periodic boundary conditions, ie imposing that $\Psi(...,x_{j}+L,...)=\Psi(...x_{j}...)$ for any $j$ yields a set of Bethe equations allowing us to determine the values of the rapidities in terms of the interaction strength and the ring circumference of length $L$:
$k_{j}=\frac{2\pi}{L}I_{j}+2\sum_{\ell}\arctan\left(\frac{k_{\ell}-k_{j}}{c}\right).$ (37)
@@
In the presence of an applied external artificial gauge field $\phi$, the many-body wavefunction has to satisfy twisted boundary conditions $\Psi(...,x_{j}+L,...)=e^{i2\pi\phi/\phi_{0}}\Psi(...x_{j}....)$. This turns onto a set of modified Bethe equations for the rapidities which read
@@ 3.2.1 Mean-field and many-body theory for fermions in the BCS-BEC crossover
We consider a two-component Fermi gas. In the case of attractive contact interactions with opposite spins, $s$-wave bound states are formed and the Fermi gas enters a superfluid phase  [^9] We will not cover here $p$-wave pairing which can occur among spin-polarized fermions with non-local interaction potentials, nor higher order type of pairing..
@@ 3.2.2 Fermi-Hubbard model
where $N_{\alpha}$ denotes the number of particles with colour $\alpha$ [^10] The appropriate units can be found by expanding $t\Delta^{2}\mathcal{H}_{FG}$, such that the first term acquires a factor of $t\Delta^{2}$ with $U\Delta$ multiplying the second. Through the effective mass $m$, we have that $t\Delta^{2}=-\frac{\hbar^{2}}{2m}$ and that $c=U\Delta/2$ thereby giving us the correspondence between the lattice and continuous parameters [286].. An appealing feature of the Gaudin-Yang-Sutherland Hamiltonian is its integrability by Bethe ansatz for all $N$, at variance with its lattice counterpart as will be discussed in Sec. 3.2.3. Consequently, through this mapping one can compare numerical results obtained for the SU($N$) Hubbard model with small filling fractions with the exact solution of the model. This allows a better understanding of the underlying system, which will be covered in more detail in Sec. 5. Note that in the presence of an artificial field, one obtains a similar Hamiltonian to Eq. (31), in that the momenta are shifted by $\frac{2\pi\phi}{L\phi_{0}}$.
@@ 3.2.3 Bethe Ansatz solution for the Fermi-Hubbard model
Sutherland discovered the exact solution of these models by generalizing the method formulated by Gaudin [292] and Yang [293] to $N$-component fermions. While for bosons, the essence of the problem lies in the exchange of momenta between two interacting particles, the situation beomes increasingly more complex when dealing with multicomponent systems having different internal states, such as SU($N$) fermions. In this case, the particles’ interaction is described through a scattering matrix that accounts for transmission and reflection, which put simply corresponds to whether an exchange of colours between the particles has occurred. To tackle the two-component case with $M$ flipped spins, Gaudin and Yang employed a clever strategy of separating the problem into the spatial and spin parts. The spatial part involves taking the $N_{p}$ particles to be indistinguishable and employing the cusp condition like the bosonic case discussed in Sec. 3.1.5. For the spin part, they realized that because of the fermions’ anti-symmetric nature, it suffices to calculate the amplitude wavefunctions for a specific set of positions of the $M$ down spins. Consequently, the problem beomes more manageable by considering a fictitious spin model describing $M$ down spins residing on a chain of length $N_{p}$ [83]. Through this observation, Yang proposed an ansatz for the spin wavefunction $\varphi(y_{1},y_{2},\ldots,y_{M})$ corresponding to the amplitudes of the inhomogenous XXX Heisenberg model, which reads
$\varphi(y_{1},y_{2},\ldots,y_{M})=\sum\limits_{\pi\in S_{M}}A(\pi)\prod\limits_{\alpha=1}^{M}F_{P}(\Lambda_{\pi\alpha;}y_{\alpha}),$ (63)
where $A(\pi)$ and $F_{P}$ are coefficients that need to be determined, $\pi$ corresponds to the set of permutations of the $M$ down spins that forms the symmetric group $S_{M}$ and $\Lambda_{\alpha}$ is the newly introduced spin rapidity analogous to the quasimomenta $k_{j}$ introduced in the spatial wavefunction. Enforcing constraints on the spin wavefunction enables one to obtain an expression for $F_{P}$ and $A(\pi)$ (see [301] for a more detailed approach). The so-called Bethe-Yang hypothesis can also be extended to multicomponent fermions, as demonstrated by Sutherland [285, 300]. Upon separating the $N_{p}$ particles into the $N$ different colours, the SU(3) case being depicted in Fig. 8, one applies the Bethe-Yang hypothesis although in a slightly different form
$\varphi(y_{1},y_{2},\ldots,y_{M})=\sum\limits_{P\in S_{M}}\psi[Q;P]\prod\limits^{M}_{n=1}F(\Lambda_{Pn},y_{Qn}),$ (64)
since now the permutation of the $M$ particles (first colour) must be considered. Subsequently, after rearranging the wavefunction coefficients $\psi[Q;P]$, the scattering matrices can be obtained, and their consistency can be checked through the Yang-Baxter relation. This procedure is repeated $N-1$ times, accounting for the extra colours in the system, with each successive application of the Bethe-Yang hypothesis seeing the introduction of a new colour $M_{r}$ having an associated spin rapidity $\Lambda^{(r)}$.
@@
for $r=1,\ldots,N-1$, with $M_{0}=N_{p}$ and $\Lambda_{\beta}^{(0)}=k_{\beta}$ with $\Phi=2\pi\phi$ [^11] In the absence of a magnetic flux, one recovers the Bethe ansatz equations corresponding to a field-free Hamiltonian. Note that the flux dependence is incorporated solely into the first equation as it is colour independent. If this were not the case, then the solvability of the model could not be guaranteed [302].. The number of particles in any given colour is given by $M_{r}$ having the associated charge and spin rapidities denoted by $k_{j}$ and $\Lambda_{\alpha}^{(r)}$. The power of the Bethe-Yang hypothesis is the reduction of the problem’s dimensionality, becoming smaller on each successive application, earning it the moniker of nested Bethe ansatz. On top of the $N_{p}$ equations as in the bosonic/one-component case, there are additional $\sum_{r}M_{r}$ equations for $N$-component systems to account for the various internal states. For the lattice model, the Bethe equations are of the same form albeit slightly different since $k_{j}\rightarrow\sin k_{j}$, which originates during the lattice regularization of the continuous system due to the different dispersion relations of the two theories [233, 156]. Considering the logarithmic form of the Bethe equations and defining $\theta_{z}(x-y)=-2\arctan(\frac{x-y}{z})$, we find
@@
which sees the introduction of the quantum numbers pertaining to the charge $I_{j}$ and spin rapidities $\{J_{\alpha},L_{a}\}$ parameterising them [^12] For the sake of convenience, we opted to denote $\Lambda_{\alpha}=\Lambda_{\alpha}^{(1)}$ and $\lambda_{a}=\Lambda_{\alpha}^{(2)}$.. Manipulation of these integer or half-odd integer numbers, mutually distinct by construction, and centered around zero for the ground-state, generates the whole spectrum ($E=-2\sum_{j}\cos(k_{j})$ or $E=\sum_{j}k_{j}^{2}$ for lattice and continuous systems respectively), which plays a pivotal role in a deeper comprehension of the system’s behaviour [301, 290].
@@ 3.2.4 Bosonizing fermionic systems
3.3 ‘Bloch theorem’ and parity dependence of persistent currents
In this section, we will discuss how the magnetic flux periodicity and patterns of persistent currents indeed arise from general properties of the system, like local particle conservation and symmetry of the wave function. To be specific, the features summarized in this section hold for a generic particle translational invariant interaction and in the presence of a static disorder.
@@
$E_{v}=E(\phi=0)+{{N\hbar^{2}}\over{2m}}\int{|\psi_{0}(\mathbf{r})|^{2}}|\nabla\chi({\mathbf{r}})|^{2},$ (75)
The next step of the argument relies on the single-valuedness of the many-body wavefunction $\psi_{0}$: if one of the particles is transported along a closed loop (keeping the other particles frozen), $\psi_{0}$ must not change. On the other hand, such an operation would involve the exchange of $N-1$ particles and because of the anti-symmetry of the wavefunction, $N-1$ sign changes. For even $N$, the single-valuedness of the wave function can hold only if $\psi_{0}$ has (at least) a further node (on top of the $N-1$ nodes coming from Pauli principle): $\psi_{0}(\mathbf{\overline{r}})=0$, where $\overline{r}$ defines a ‘non-symmetry dictated nodal surface’. The phase $\chi({\mathbf{r}})$ can be chosen in the form of a phase twist, namely with a phase discontinuity concentrated at $\overline{r}$. This choice of $\chi({\mathbf{r}})$ can be made without resulting in an increase in energy (due to the aforementioned property $\psi_{0}(\mathbf{\overline{r}})=0$). In this case, the variational energy results to be $E_{v}=E(\phi=0)$ and then an upper bound of the exact energy $E(\phi)$: $E(\phi)\geq E(\phi=0)$. Following a similar argument, it can be concluded that, for odd $N$, $E(\phi)\geq E(\phi=\pi)$. By resorting to specific inequalities fulfilled by the path integral representation of the partition function of the system, lower bounds to the many-body energy were obtained by Waintal et al. [161]. Combining the Leggett and Waintal results, the ground state energy is constrained as
$E(\phi=0)\leq E(\phi)\leq E(\phi=\pi)\,,\quad N\,odd,$ (76)
$E(\phi=\pi)\leq E(\phi)\leq E(\phi=0)\,,\quad N\,even\,.$ (77)
@@
$F(\phi=0)\leq E(\phi)\,,\quad N_{\uparrow},N_{\downarrow}\quad odd,$ (78)
$F(\phi=\pi)\leq E(\phi)\,,\quad N_{\uparrow},N_{\downarrow}\quad even.$ (79)

### v3

@@ 3.1.2 The Lieb-Liniger one dimensional model for interacting bosons
with $x_{j}$ being the position of the $j$-th particle on the ring and $\phi_{0}$ corresponding to the bare flux quantum. The solution of this model provides us with a full many-body tractable model of interacting bosons in a finite-size ring that can sustain persistent currents. The Lieb-Liniger model is integrable through Bethe-ansatz [240] and will be discussed in detail in the following section. Its integrability and exact solution allow us to use this model as a test bench and reference model to understand the microscopic origin of the phenomena observed in more complex systems. For this reason, a plethora of works have and still do investigate bosonic systems using the Lieb-Liniger model [241, 242, 243, 125]. Despite its integrability through Bethe Ansatz, calculating high-order correlations using the exact solution becomes difficult. New techniques have been proposed, for example the use of quantum computers, in which one can prepare the quantum state to later perform measures on it [244, 245, 246].
The two models presented, GPE and LL, describe bosonic interacting particles in the continuum. However we remark again that the first is based on mean-field theory and therefore restricted to the limit of weak interactions and a large number of particles, while the latter describes the full-many body wavefunction exactly. Both theories can attain attractive, $a_{s}<0$, and repulsive $a_{s}>0$ interactions. For repulsive interactions, both models give a ground state with a homogeneous density along the ring because of the translational invariance present in the system. Intuitively, repulsion tends to spread particles apart, and therefore, a homogenous density keeps them as separated as possible without creating regions of higher density that would lead to increased local interactions. On the other hand, for attractive interactions, one could naively expect the localization of particles in space due to the negative two-body contact interactions. Such localization would, in principle, break translational invariance, and it is found in the mean-field description for sufficiently strong attractive interactions [247]. This mean-field states were proposed as candidates to present time-crystal-like behavior and later disproved [248, 249]. However, we want to clarify that the apparent breaking of translational invariance is only observed in single-shot density measurements where the particle’s center-of-mass is localized at a given position. One can also access such localization by calculating expectation values of higher order correlators. We remark that, as pointed out in [241], the alleged breaking of translational invariance is because one would need to average over the center-of-mass position of the mean-field solution in order to recover the correct translational invariant solution. This connection between mean-field and many-body theories were also investigated in Ref. [250]. Thus, here we speak of current states in the attractive regime, we consider the ground state to be a delocalized state of bound-state particles.
@@ 3.1.3 The Bose-Hubbard model
where $b^{\dagger}_{j}$ ($b_{j}$) creates (annihilates) a boson, and $n_{j}=b_{j}^{\dagger}b_{j}$ is the local particle number operator for site $j$. The parameters $J$ and $U$ account for the hopping amplitude and on-site interaction when restricting the particles to occupy the lowest Bloch band in each lattice site respectively [251]. The elementary flux quantum is denoted by $\phi_{0}$ and is taken to be 1 unless explicitly stated. The BHM has been experimentally realized in experiments [67, 69] with remarkable control on the parameters of the systems and read-out of the particles trapped in the lattice. With such a model, they probe the superfluid to Mott insulator transition in a highly controllable environment. This model also provides a full many-body description of the interacting bosons. However, due to its imposed restriction provided by the lattice, the Hilbert space dimensions become approachable from a numerical perspective, with techniques such as Exact Diagonalization or density matrix renormalization group (DMRG) [252, 253].
@@ 3.1.5 Exact solutions for interacting bosons in 1D: Bethe Ansatz and Tonks-Girardeau
Repeated use of Eq. (33) together with the imposition of the periodic boundary conditions, ie imposing that $\Psi(...,x_{j}+L,...)=\Psi(...x_{j}...)$ for any $j$ yields a set of Bethe equations allowing us to determine the values of the rapidities in terms of the interaction strength and the ring circumference of length $L$. In the presence of an applied external artificial gauge field $\phi$, the many-body wavefunction has to satisfy twisted boundary conditions $\Psi(...,x_{j}+L,...)=e^{i2\pi\phi/\phi_{0}}\Psi(...,x_{j},....)$. This turns onto a set of modified Bethe equations for the rapidities which read
@@ 3.2.1 Mean-field and many-body theory for fermions in the BCS-BEC crossover
We consider a two-component Fermi gas. In the case of attractive contact interactions with opposite spins, $s$-wave bound states are formed and the Fermi gas enters a superfluid phase.(6)(6)(6)We will not cover here $p$-wave pairing which can occur among spin-polarized fermions with non-local interaction potentials, nor higher order type of pairing.
@@ 3.2.2 Fermi-Hubbard model
where $N_{\alpha}$ denotes the number of particles with colour(7)(7)(7)The appropriate units can be found by expanding $t\Delta^{2}\mathcal{H}_{FG}$, such that the first term acquires a factor of $t\Delta^{2}$ with $U\Delta$ multiplying the second. Through the effective mass $m$, we have that $t\Delta^{2}=-\frac{\hbar^{2}}{2m}$ and that $c=U\Delta/2$ thereby giving us the correspondence between the lattice and continuous parameters [308]. $\alpha$. An appealing feature of the Gaudin-Yang-Sutherland Hamiltonian is its integrability by Bethe ansatz for all $N$, at variance with its lattice counterpart as will be discussed in Sec. 3.2.3. Consequently, through this mapping one can compare numerical results obtained for the SU($N$) Hubbard model with small filling fractions with the exact solution of the model. This allows a better understanding of the underlying system, which will be covered in more detail in Sec. 5. Note that in the presence of an artificial field, one obtains a similar Hamiltonian to Eq. (28), in that the momenta are shifted by $\frac{2\pi\phi}{L\phi_{0}}$.
@@ 3.2.3 Bethe Ansatz solution for the Fermi-Hubbard model
Sutherland discovered the exact solution of these models by generalizing the method formulated by Gaudin [314] and Yang [315] to $N$-component fermions. While for bosons, the essence of the problem lies in the exchange of momenta between two interacting particles, the situation becomes increasingly more complex when dealing with multicomponent systems having different internal states, such as SU($N$) fermions. In this case, the particles’ interaction is described through a scattering matrix that accounts for transmission and reflection, which simply put corresponds to whether an exchange of colours between the particles has occurred. To tackle the two-component case with $M$ flipped spins, Gaudin and Yang employed a clever strategy of separating the problem into the spatial and spin parts. The spatial part involves taking the $N_{p}$ particles to be indistinguishable and employing the cusp condition like the bosonic case discussed in Sec. 3.1.5. For the spin part, they realized that because of the fermions’ anti-symmetric nature, it suffices to calculate the amplitude wavefunctions for a specific set of positions of the $M$ down spins. Consequently, the problem becomes more manageable by considering a fictitious spin model describing $M$ down spins residing on a chain of length $N_{p}$ [86]. Through this observation, Yang proposed an ansatz for the spin wavefunction $\varphi(y_{1},y_{2},\ldots,y_{M})$ corresponding to the amplitudes of the inhomogenous XXX Heisenberg model, with a newly introduced spin rapidity $\Lambda_{\alpha}$ analogous to the quasimomenta $k_{j}$ introduced in the spatial wavefunction (see [323, 175, 312] for a more detailed approach). This so-called Bethe-Yang hypothesis can also be extended to multicomponent fermions, as demonstrated by Sutherland [307, 322]. Upon separating the $N_{p}$ particles into the $N$ different colours, the SU(3) case being depicted in Fig. 8, one applies the Bethe-Yang hypothesis although in a slightly different form (see [307]). Repeating the nested ansatz $N-1$ times, to account for the extra colours in the system, each successive application of the Bethe-Yang hypothesis sees the introduction of a new colour $M_{r}$ having an associated spin rapidity $\Lambda^{(r)}$.
@@
for $r=1,\ldots,N-1$, with $M_{0}=N_{p}$ and $\Lambda_{\beta}^{(0)}=k_{\beta}$ with(8)(8)(8)In the absence of a magnetic flux, one recovers the Bethe ansatz equations corresponding to a field-free Hamiltonian. Note that the flux dependence is incorporated solely into the first equation as it is colour independent. If this were not the case, then the solvability of the model could not be guaranteed [324]. $\Phi=2\pi\phi$. The number of particles in any given colour is given by $M_{r}$ having the associated charge and spin rapidities denoted by $k_{j}$ and $\Lambda_{\alpha}^{(r)}$. The power of the Bethe-Yang hypothesis is the reduction of the problem’s dimensionality, becoming smaller on each successive application, earning it the moniker of nested Bethe ansatz. On top of the $N_{p}$ equations as in the bosonic/one-component case, there are additional $\sum_{r}M_{r}$ equations for $N$-component systems to account for the various internal states. For the lattice model, the Bethe equations are of the same form albeit slightly different since $k_{j}\rightarrow\sin k_{j}$, which originates during the lattice regularization of the continuous system due to the different dispersion relations of the two theories [255, 175]. Considering the logarithmic form of the Bethe equations and defining $\theta_{z}(x-y)=-2\arctan(\frac{x-y}{z})$, we find
@@
which sees the introduction of the quantum numbers pertaining to the charge $I_{j}$ and spin rapidities $\{J_{\alpha},L_{a}\}$ parameterising them.(9)(9)(9)For the sake of convenience, we opted to denote $\Lambda_{\alpha}=\Lambda_{\alpha}^{(1)}$ and $\lambda_{a}=\Lambda_{\alpha}^{(2)}$. Manipulation of these integer or half-odd integer numbers, mutually distinct by construction, and centered around zero for the ground-state, generates the whole spectrum ($E=-2\sum_{j}\cos(k_{j})$ or $E=\sum_{j}k_{j}^{2}$ for lattice and continuous systems respectively), which plays a pivotal role in a deeper comprehension of the system’s behaviour [323, 312].
@@ 3.2.4 Bosonizing fermionic systems
3.3 ‘Bloch bands’ and parity dependence of persistent currents
In this section, we will discuss how the magnetic flux periodicity and patterns of persistent currents indeed arise from the general properties of the system, like local particle conservation and symmetry of the wave function. To be specific, the features summarized in this section hold for a generic particle translational invariant interaction and in the presence of a static disorder.
@@
$E_{v}=E(\phi=0)+{{N_{p}\hbar^{2}}\over{2m}}\int{|\psi_{0}(\mathbf{r})|^{2}}|\nabla\chi({\mathbf{r}})|^{2},$ (69)
The next step of the argument relies on the single-valuedness of the many-body wavefunction $\psi_{0}$: if one of the particles is transported along a closed loop (keeping the other particles frozen), $\psi_{0}$ must not change. On the other hand, such an operation would involve the exchange of $N_{p}-1$ particles and because of the anti-symmetry of the wavefunction, $N_{p}-1$ sign changes. For even $N_{p}$, the single-valuedness of the wave function can hold only if $\psi_{0}$ has (at least) a further node (on top of the $N_{p}-1$ nodes coming from Pauli principle): $\psi_{0}(\mathbf{\overline{r}})=0$, where $\overline{r}$ defines a ‘non-symmetry dictated nodal surface’. The phase $\chi({\mathbf{r}})$ can be chosen in the form of a phase twist, namely with a phase discontinuity concentrated at $\overline{r}$. This choice of $\chi({\mathbf{r}})$ can be made without resulting in an increase in energy (due to the aforementioned property $\psi_{0}(\mathbf{\overline{r}})=0$). In this case, the variational energy results to be $E_{v}=E(\phi=0)$ and then an upper bound of the exact energy $E(\phi)$: $E(\phi=0)\geq E(\phi)$. Following a similar argument, it can be concluded that, for odd $N_{p}$, $E(\phi=\pi)\geq E(\phi)$. Eventually, Loss proved this conjecture in the Luttinger liquid theory framework through the bosonization method [329]. By resorting to specific inequalities fulfilled by the path integral representation of the partition function of the system, lower bounds to the many-body energy were obtained by Waintal et al. [331]. Combining the Leggett and Waintal results, the ground state energy is constrained as
$E(\phi=0)\leq E(\phi)\leq E(\phi=\pi)\,,\quad N_{p}\,odd,$ (70)
$E(\phi=\pi)\leq E(\phi)\leq E(\phi=0)\,,\quad N_{p}\,even\,.$ (71)
@@
$F(\phi=0)\leq F(\phi)\,,\quad N_{p\uparrow},N_{p\downarrow}\quad odd,$ (72)
$F(\phi=\pi)\leq F(\phi)\,,\quad N_{p\uparrow},N_{p\downarrow}\quad even.$ (73)
These parity effects also affects the fermionic current as it causes parabolas of different angular momenta $\ell$ to traverse each other at $\phi=\frac{j}{2}\phi_{0}$ or $\phi=j\phi_{0}$ for integer $j$ depending on whether $N_{p}$ is odd or even respectively. Essentially, the landscape of the current for spinless fermions indicates that they act as a diamagnet (paramagnet) for odd (even) $N_{p}$ [173] in the presence of an effective magnetic flux. For electrons/two-component fermions, i.e. having a spin-up and spin-down, this amounts to saying that the current has a diamagnetic (paramagnetic) character flow in a system consisting of $N_{p}=4\nu+2$ ($N_{p}=4\nu$) particles for $\nu\in\mathbb{Z}^{+}$ and equal population per component [332, 331]. Systems with a larger number of components will be addressed in Sec. 5.3. For bosons, there is no parity effect as the current is always diamagnetic due to the lack of Pauli exclusion principle. Another noteworthy feature is that the circulation of the particles, be they moving clockwise or anti-clockwise, is captured through the orientation of the persistent current on account of its parity.